# Shifting AI Efficiency From Model-Centric to Data-Centric Compression

### v2

@@ 3.1 Research Roadmap - What Makes Token Compression Work?
Existing token compression methods fundamentally operate through a two-stage process: first, identifying tokens eligible for compression within the existing token sequence $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}]$ using carefully designed compression criteria through a scoring function $\mathcal{E}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, and then determining the precise handling of these tokens through specific compression strategies $\mathcal{P}:(\mathbf{X},\{s_{t}\}_{t=1}^{T})\to\mathbf{X}^{\prime}$ that transform the original sequence into a compressed one where $|\mathbf{X}^{\prime}|<|\mathbf{X}|$. Given that existing research primarily revolves around these two key components, we next systematically analyze their designs and review representative approaches across various scenarios.
@@
(I) Parametric Methods employ auxiliary networks as scoring functions $\mathcal{E}_{\Delta\theta}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, introducing additional parameters $\Delta\theta$ beyond the original model parameters $\theta$. These methods include: (i) training-aware approaches Rao2021:DynamicViT ; you2024DiffCR ; li2024tokenpacker ; kim2022learned that optimize $\Delta\theta$ through training to learn scoring function $\mathcal{S}_{\Delta\theta}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, and (ii) training-free approaches Mahmud2024:PaPr ; zhao2024SGL that directly employ pre-trained networks as scoring function $\mathcal{S}_{\text{fixed}}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$ without updating $\Delta\theta$.
(II) Non-parametric Methods utilize parameter-free heuristics for token scoring without introducing extra parameters. These approaches can be categorized into: (i) inherent computation methods Liang2022:EViT ; zou2025:ToCa ; Chen2024:FastV ; xiao2024efficient ; ge2024model that leverage model’s internal calculations for token scoring $\mathcal{S}_{\text{in}}:\mathbf{A}\to\{s_{t}\}_{t=1}^{T}$, such as using attention weights ($s_{t}=\sum_{j=1}^{T}a_{t}^{j}$, where $a_{t}^{j}$ represents attention score between tokens), and (ii) external computation methods bolya2023tome ; Zhang2025:SiTo ; devoto-etal-2024-simple ; wang2024model ; liu2025vidcom2 that design additional metrics $\mathcal{S}_{\text{ex}}:\mathbf{Z}\to\mathbb{R}^{T\times T}$ to evaluate token relationships. For external methods, an additional function $g:\mathbf{X}\to\mathbf{Z}$ is introduced to compute intermediate features, where $\mathbf{Z}=g(\mathbf{X})$. The scoring function then operates on these features: $s_{i,j}=f(\mathbf{z}_{i},\mathbf{z}_{j})$, where $f$ is a custom pairwise scoring function. A typical example is using cosine similarity, where $g$ is an identity function and $s_{i,j}=\frac{\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle}{\|\mathbf{x}_{i}\|_{2}\|\mathbf{x}_{j}\|_{2}}$.
@@
To reduce sequence length while preserving critical information, compression strategies $\mathcal{P}$ transform the original sequence based on token scores $\{s_{t}\}_{t=1}^{T}$. These strategies can be primarily categorized into two distinct approaches:
(I) Token Pruning directly discards less important tokens from the sequence based on their scores. These methods Rao2021:DynamicViT ; goyal2020power ; jiang2023llmlingua ; Chen2024:FastV typically remove tokens with scores below a threshold, producing a compressed sequence:
@@
where $\tau$ is a threshold determining token removal. Token pruning reduces computation through direct elimination but risks information loss, particularly for fine-grained tasks.
(II) Token Merging preserves information by combining semantically similar tokens bolya2023tome ; Liang2022:EViT ; zhang2024sparsevlm ; Bolya2023:ToMeSD . Given an input sequence $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}$ and a mapping $\pi:\{1,\dots,T\}\to\{1,\dots,M\}$ that assigns tokens to $M$ merge groups based on their semantic relationships, this approach generates a compressed sequence $\mathbf{X}^{\prime}=\{\mathbf{x}^{\prime}_{1},\dots,\mathbf{x}^{\prime}_{M}\}$ through weighted aggregation:
@@
3.2 Training and Inference Targets - How Token Compression Benefits?
@@
Token compression techniques contribute to improving both the quality and efficiency of model training. These benefits can be broadly categorized into two aspects: enhancing training quality and increasing training efficiency.
@@
$\bullet$ Data augmentation techniques have been widely adopted to enrich training datasets by introducing variability that enhances robustness and informativeness Autoaugment . In computer vision, mixing or combining image tokens creates novel representations that elevate training effectiveness mixup ; cutmix . This strategy has also been extended to synthetic datasets, where adaptive augmentation controls the informativeness of generated images DCC ; DSA ; NCFM . Analogously, in natural language processing, augmenting text tokens through synonym replacement wei2019eda , contraction expansion coulombe2018text , back-translation chen2020mixtext , and reformulation hao2025maga , supporting better generalization.
$\bullet$ Token selection focuses on filtering out low-quality tokens to refine training data quality rho1 ; lee2022deduplicating ; penedo2023refinedweb ; wenzek2019ccnet ; gao2020pile ; li2024datacomp . Common approaches include rule-based heuristics raffel2020exploring ; penedo2024fineweb , deduplication methods lee2022deduplicating ; penedo2023refinedweb ; abbas2023semdedup , and scoring strategies leveraging large language models wenzek2019ccnet ; gao2020pile ; li2024datacomp ; wettig2024qurating ; sachdeva2024train .
Formally, consider a training batch =$\mathcal{B}=\{\mathbf{X}_{i}\}_{i=1}^{N}$, where each $\mathbf{X}_{i}=[\mathbf{x}_{i,1},\mathbf{x}_{i,2},\dots,\mathbf{x}_{i,T}]$ is a token sequence of length $T$. A quality scoring function $q:\mathcal{T}\to\mathbb{R}$ assigns each token $\mathbf{x}_{i,j}\in\mathcal{T}$ a score reflecting its informativeness or relevance. Using a threshold $\tau$, tokens with scores below $\tau$ are filtered out via a mask $\mathbf{m}_{i}$: $m_{i,j}=\begin{cases}1,&q(\mathbf{x}_{i,j})\geq\tau\\ 0,&\text{otherwise}\end{cases}$. The filtered batch $\tilde{\mathcal{B}}$ consists of sequences:
@@
Training on these curated, high-quality tokens enables the model to concentrate on the most relevant information, reducing noise and redundancy, which enhances generalization and learning efficiency.
(II) Increasing Training Efficiency Token compression directly reduces the token length processed during training, addressing critical challenges associated with scaling large models Bolya2023:ToMe ; Choudhury2024:RLT ; Shang2024:LLaVA-PruMerge ; xing2024pyramiddrop . For Transformer architectures with sequence length reduced from $n$ to $m$ ($m<n$), the computational and memory benefits can be quantified as:
$\frac{\Omega(\mathbf{X}^{\prime})}{\Omega(\mathbf{X})}=\frac{\mathcal{O}(m^{2}d)}{\mathcal{O}(n^{2}d)}=\mathcal{O}\left(\frac{m^{2}}{n^{2}}\right),\quad\frac{\mathcal{M}(\mathbf{X}^{\prime})}{\mathcal{M}(\mathbf{X})}\approx\frac{md}{nd}=\frac{m}{n},$ (7)
where $d$ is the embedding dimension, $\Omega(\cdot)$ represents the computational measure, and $\mathcal{M}(\cdot)$ denotes the memory measure. This quadratic reduction in computation and linear reduction in memory enable faster training iterations and larger batch sizes on fixed hardware resources.
@@
Token compression methods can also enhance model inference efficiency through two key aspects: decreasing computational complexity and reducing memory usage.
(I) Decreasing Computational Complexity Following patterns established in training, token compression achieves quadratic speedup in inference computations. Notably, many non-parametric compression methods Bolya2023:ToMe ; Ju2024:Turbo ; Chen2024:FastV can be directly integrated into inference without additional training or architectural modifications, enabling immediate benefits across various domains zhang2024sparsevlm ; wen2025stop .
(II) Reducing Memory Usage Token compression optimizes memory efficiency through two mechanisms: (i) computing memory reduction following the linear scaling pattern shown in training, and (ii) KV cache optimization for large language models li2024snapkv ; cai2024pyramidkv ; wan2024look-m ; wan2025meda . During autoregressive generation, each layer caches key and value states for attention computation, with memory growing with sequence length. For a sequence of length $n$ compressed to length $m$, with $L$ layers and hidden dimension $d$, the KV cache memory reduction is:
@@
These benefits are particularly crucial for real-time interactive systems, including UI agents tang2025guisurvey ; tang2025gui , autonomous driving gao2021autonomous , and embodied AI duan2022embodied ; yang2025efficientvla , where efficient processing of continuous inputs under resource constraints is essential.
3.3 Compelling Advantages - Why Token Compression Matters?
Based on our comprehensive analysis of token compression techniques, we identify five compelling advantages that make them particularly promising:
  • Universal Applicability: The redundancy of tokens exists consistently across modalities and tasks, making token compression possible in all kinds of settings.
  • Dual-phase Efficiency: Token compression is capable of accelerating both model training and inference phases with minimal accuracy loss.
  • Architectural Compatibility: Token compression is orthogonal to existing model compression and compression methods, making it is possible to be integrated seamlessly with existing compression techniques. Besides, it is friendly to the hardware and computer systems.
  • Low Implementation Costs: Modern neural networks, such as transformers, is able to process tokens of different lengths. As a result, token compression can be done without introducing any training costs and data utilization.
  • Quadratic Gains: The $\mathcal{O}(n^{2})$ computation complexity of widely used self-attention indicates token compression can bring significant benefits in computation.
As AI development enters a new phase where context length becomes the primary bottleneck, the research focus of AI efficiency should shift towards data-centric compression through token compression, enabling more efficient and scalable AI systems.
@@ 4.1 Performance Degradation
Methodological Bottlenecks. Attention scores play a crucial role across existing token compression approaches. For example, [CLS] token attention scores are used to select key visual tokens qiang2022attcat ; haurum2023tokens ; zhang2024cls ; wang2024cls ; han2024rethinking ; yang2024visionzip ; liu2025compression , while cross-modal guidance jiang2022trips ; cao2023pumer ; shi2023crossget ; chen2024image ; xing2024pyramiddrop ; zhang2024sparsevlm relies on text-vision attention scores. But are attention scores truly reliable for deciding which tokens to keep? Recent work zhang2024cls ; wen2025token reveals that attention scores can suffer from position bias. For instance, when using text-vision scores to retain visual tokens, those near the sequence end often get higher weights. In 2D image space, this biases retention toward the lower half or bottom-right corner. Clearly, it’s unrealistic to assume the lower half of all images is more important. Such bias can significantly hurt compression performance. As shown in Figure 2, a comprehensive review and comparison of key studies jiang2025kind ; wen2025stop ; wen2025token ; Zhang2024:SparseVLM ; liu2024multi ; liu2025vidcom2 confirms our hypothesis: even well-crafted attention-based methods can underperform simple random pruning or pooling. Detailed analysis is in Appendix B.
@@
Inherent Limitations of Token Compression. Beyond performance degradation from methodological design, does token compression face inherent limitations? Is it universally applicable across various tasks? For multi-modal large language models, wen2025token shows most existing compression methods underperform on visual grounding tasks, with significant drops on benchmarks like RefCOCO yu2016modeling . In OCR-related parsing yang2024cc ; ouyang2024omnidocbench , documents with dense layouts yield highly information-rich visual tokens. Compressing these risks severe information loss and degraded performance. Beyond vision, current methods also face inherent limits in other modalities. In Automatic Speech Recognition (ASR) and Automatic Speech Translation (AST) ardila2019common ; conneau2023fleurs , audio is encoded and then decoded into text using an MLLM abouelenin2025phi ; chu2024qwen2 . Audio tokens in ASR and AST are dense and temporally continuous. Pruning or merging them disrupts this continuity, leading to fragmented recognition or translation. Similarly, translation across languages in the text modality may suffer significant degradation under high compression ratios.
4.2 Suboptimal Token Representation

### v3

@@ 3.1 Research Roadmap - What Makes Data-centric Compression Work?
Existing data-centric compression methods (i.e., token compression and dataset compression) fundamentally operate through a two-stage process (see Figure 2): first, identifying tokens eligible for compression within the existing token sequence $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}]$ using carefully designed compression criteria through a scoring function $\mathcal{E}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, and then determining the precise handling of these tokens through specific compression strategies $\mathcal{P}:(\mathbf{X},\{s_{t}\}_{t=1}^{T})\to\mathbf{X}^{\prime}$ that transform the original sequence into a compressed one where $|\mathbf{X}^{\prime}|<|\mathbf{X}|$. Given that existing research primarily revolves around these two key components, we next systematically analyze their designs and review representative approaches.
Figure 2: Overview of the data-centric compression paradigm. Given an input token sequence $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{T}]$, data-centric compression first computes importance scores via a scoring function $\mathcal{E}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, then generates a compressed sequence $\mathbf{X}^{\prime}$ through a compression strategy $\mathcal{P}:(\mathbf{X},\{s_{t}\}_{t=1}^{T})\to\mathbf{X}^{\prime}$, where $|\mathbf{X}^{\prime}|<|\mathbf{X}|$.
@@
(I) Parametric Methods employ auxiliary networks as scoring functions $\mathcal{E}_{\Delta\theta}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, introducing additional parameters $\Delta\theta$ beyond the original model parameters $\theta$. These methods include: (i) training-aware approaches Rao et al. (2021); You et al. (2024); Li et al. (2024d); Kim et al. (2022) that optimize $\Delta\theta$ through training to learn scoring function $\mathcal{S}_{\Delta\theta}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, and (ii) training-free approaches Mahmud et al. (2024); Zhao et al. (2024a) that directly employ pre-trained networks as scoring function $\mathcal{S}_{\text{fixed}}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$ without updating $\Delta\theta$.
(II) Non-parametric Methods utilize parameter-free heuristics for token scoring without introducing extra parameters. These approaches can be categorized into: (i) inherent computation methods Liang et al. (2022); Zou et al. (2025); Chen et al. (2024b); Xiao et al. (2024); Ge et al. (2024) that leverage model’s internal calculations for token scoring $\mathcal{S}_{\text{in}}:\mathbf{A}\to\{s_{t}\}_{t=1}^{T}$, such as using attention weights ($s_{t}=\sum_{j=1}^{T}a_{t}^{j}$, where $a_{t}^{j}$ represents attention score between tokens), and (ii) external computation methods Bolya et al. (2023a); Zhang et al. (2025a); Devoto et al. (2024); Wang et al. (2024c); Liu et al. (2025a) that design additional metrics $\mathcal{S}_{\text{ex}}:\mathbf{Z}\to\mathbb{R}^{T\times T}$ to evaluate token relationships. For external methods, an additional function $g:\mathbf{X}\to\mathbf{Z}$ is introduced to compute intermediate features, where $\mathbf{Z}=g(\mathbf{X})$. The scoring function then operates on these features: $s_{i,j}=f(\mathbf{z}_{i},\mathbf{z}_{j})$, where $f$ is a custom pairwise scoring function. A typical example is using cosine similarity, where $g$ is an identity function and $s_{i,j}=\frac{\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle}{\|\mathbf{x}_{i}\|_{2}\|\mathbf{x}_{j}\|_{2}}$.
@@
To reduce sequence length while preserving critical information, compression strategies $\mathcal{P}$ transform the original sequence based on token scores $\{s_{t}\}_{t=1}^{T}$. These strategies can be categorized into two approaches:
(I) Token Pruning directly discards less important tokens from the sequence based on their scores. These methods Rao et al. (2021); Goyal et al. (2020); Jiang et al. (2023); Chen et al. (2024b) typically remove tokens with scores below a threshold, producing a compressed sequence:
@@
where $\tau$ is a threshold determining token removal. Token pruning reduces computation through direct elimination but risks information loss, particularly for fine-grained tasks Xie et al. (2021).
(II) Token Merging preserves information by combining semantically similar tokens Bolya et al. (2023a); Zhang et al. (2025b); Bolya and Hoffman (2023). Given an input sequence $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}$ and a mapping $\pi:\{1,\dots,T\}\to\{1,\dots,M\}$ that assigns tokens to $M$ merge groups based on their semantic relationships, this approach generates a compressed sequence $\mathbf{X}^{\prime}=\{\mathbf{x}^{\prime}_{1},\dots,\mathbf{x}^{\prime}_{M}\}$ through weighted aggregation:
@@
3.2 Training and Inference Targets - How Data-centric Compression Benefits?
@@
Data-centric compression methods contribute to improving both the quality and efficiency of model training. These benefits can be broadly categorized into two aspects: enhancing training quality and increasing training efficiency.
@@
(i) Data augmentation techniques have been widely adopted to enrich training datasets by introducing variability that enhances robustness and informativeness Cubuk et al. (2018). In computer vision, mixing or combining image tokens creates novel representations that elevate training effectiveness Zhang et al. (2018); Yun et al. (2019). This strategy has also been extended to synthetic datasets, where adaptive augmentation controls the informativeness of generated images Zhao and Bilen (2021); Lee et al. (2022b); Wang et al. (2025d, c, 2024b). Analogously, in natural language processing, augmenting text tokens through synonym replacement Wei and Zou (2019), contraction expansion Coulombe (2018), back-translation Chen et al. (2020), and reformulation Hao et al. (2025), supporting better generalization.
(ii) Token selection focuses on filtering out low-quality tokens to refine training data quality Lin et al. (2024b); Lee et al. (2022a); Penedo et al. (2023); Wenzek et al. (2019); Gao et al. (2020); Li et al. (2024b); Wang et al. (2025b). Common approaches include rule-based heuristics Raffel et al. (2020); Penedo et al. (2024), deduplication methods Lee et al. (2022a); Penedo et al. (2023); Abbas et al. (2023), and scoring strategies leveraging large language models Wenzek et al. (2019); Gao et al. (2020); Li et al. (2024b); Wettig et al. (2024); Sachdeva et al. (2024); Wang et al. (2025a).
Formally, consider a training batch $\mathcal{B}=\{\mathbf{X}_{i}\}_{i=1}^{N}$, where each $\mathbf{X}_{i}=[\mathbf{x}_{i,1},\mathbf{x}_{i,2},\dots,\mathbf{x}_{i,T}]$ is a token sequence of length $T$. A quality scoring function $q:\mathcal{T}\to\mathbb{R}$ assigns each token $\mathbf{x}_{i,j}\in\mathcal{T}$ a score reflecting its informativeness or relevance. Using a threshold $\tau$, tokens with scores below $\tau$ are filtered out via a mask $\mathbf{m}_{i}$: $m_{i,j}=\begin{cases}1,&q(\mathbf{x}_{i,j})\geq\tau\\ 0,&\text{otherwise}\end{cases}$. The filtered batch $\tilde{\mathcal{B}}$ consists of sequences:
@@
Training on these curated, high-quality tokens enables the model to focus on the most relevant information, reducing noise and redundancy, thereby improving generalization and learning efficiency.
(II) Increasing Training Efficiency Data-centric compression methods directly reduce the sequence length processed during training, addressing critical challenges associated with scaling large models Bolya et al. (2023b); Choudhury et al. (2024); Shang et al. (2024); Xing et al. (2025). For Transformer architectures with sequence length reduced from $n$ to $m$ ($m<n$), the computational and memory benefits can be quantified as:
$\frac{\Omega(\mathbf{X}^{\prime})}{\Omega(\mathbf{X})}$ $=\frac{\mathcal{O}(m^{2}d)}{\mathcal{O}(n^{2}d)}=\mathcal{O}\left(\frac{m^{2}}{n^{2}}\right),$ (7)
$\frac{\mathcal{M}(\mathbf{X}^{\prime})}{\mathcal{M}(\mathbf{X})}$ $\approx\frac{md}{nd}=\frac{m}{n}.$
where $d$ is the embedding dimension, $\Omega(\cdot)$ represents the computational measure, and $\mathcal{M}(\cdot)$ denotes the memory measure. This quadratic reduction in computation and linear reduction in memory enables faster training iterations and larger batch sizes on fixed hardware resources.
@@
Data-centric compression methods can also enhance model inference efficiency through two key aspects: decreasing computational complexity and reducing memory usage.
(I) Decreasing Computational Complexity: Following patterns established in training, data-centric compression achieves quadratic speedup in inference computations. Notably, many non-parametric compression methods Bolya et al. (2023b); Chen et al. (2024b) can be directly integrated into inference without additional training or architectural modifications, enabling immediate benefits across domains Zhang et al. (2025b); Wen et al. (2025b).
(II) Reducing Memory Usage: Data-centric compression optimizes memory efficiency through two mechanisms: (i) computing memory reduction following the linear scaling pattern shown in training, and (ii) KV cache optimization for large language models Li et al. (2024e); Cai et al. (2024); Wan et al. (2024, 2025b). During autoregressive generation, each layer caches key and value states for attention computation, with memory growing with sequence length. For a sequence of length $n$ compressed to length $m$, with $L$ layers and hidden dimension $d$, the KV cache memory reduction is:
@@
These benefits are particularly crucial for real-time interactive systems, including UI agents Tang et al. (2025b, a), autonomous driving Gao et al. (2021), and embodied AI Duan et al. (2022); Yang et al. (2025c), where efficient processing of continuous inputs under resource constraints is essential.
3.3 Compelling Advantages - Why Data-centric Compression Matters?
Based on comprehensive analysis of data-centric compression, we identify five compelling advantages that makes them particularly promising:
  1. 1.
    Universal Applicability: Token redundancy is consistent across modalities and tasks, enabling data-centric compression in diverse settings.
  2. 2.
    Dual-phase Efficiency: Data-centric compression accelerates both training and inference with minimal accuracy loss.
  3. 3.
    Architectural Compatibility: Data-centric compression is orthogonal to compression methods and integrates seamlessly with them. It is also hardware and system friendly.
  4. 4.
    Low Implementation Costs: Modern architectures like transformers support variable-length inputs, allowing data-centric compression without retraining or data overhead.
  5. 5.
    Quadratic Gains: The $\mathcal{O}(n^{2})$ complexity of self-attention ensures data-centric compression yields substantial computational savings.
As AI development enters a new phase where context length becomes the primary bottleneck, the research focus of AI efficiency should shift towards data-centric compression, enabling more efficient and scalable AI systems.
@@ 4.1 Performance Degradation
Figure 3: Empirical comparison of carefully designed data-centric compression methods and random token dropping. Results demonstrate that in multiple scenarios (e.g., LLMs, MLLMs, and DiTs), some carefully designed methods surprisingly underperform compared to random token selection.
@@
Methodological Bottlenecks. Attention scores are central to most data-centric compression approaches. For example, [CLS] token attention scores are used to select key visual tokens Haurum et al. (2023); Zhang et al. (2024c); Han et al. (2024b); Yang et al. (2024a); Liu et al. (2025b), while cross-modal guidance Chen et al. (2024c); Xing et al. (2025); Zhang et al. (2025b) relies on text-vision attention scores. But are attention scores truly reliable for deciding which data to retain? Recent work Zhang et al. (2024c); Wen et al. (2025a) reveals that attention scores can suffer from position bias. For instance, when using text-vision scores in LLMs to retain visual tokens, those near the sequence end often receive higher weights. In 2D image space, this biases retention toward the lower half or bottom-right corner. Clearly, it is unrealistic to assume the lower half of all images is universally more important. Such bias can significantly hurt compression performance. In Figure 3, this phenomenon is consistently observed across multiple tasks and models. Recent studies Jiang et al. (2025); Wen et al. (2025b, a); Liu et al. (2025a) have also confirmed that even well-crafted attention-based methods may underperform simple random dropping. Detailed analysis is in Appendix B.
@@
Inherent Limitations of Data-Centric Compression. Beyond methodological design, does data-centric compression face inherent limitations? Is it universally applicable across tasks? For MLLMs, Wen et al. shows most existing methods underperform on visual grounding tasks, with significant drops on benchmarks like RefCOCO Yu et al. (2016). In OCR-related parsing Yang et al. (2024b); Ouyang et al. (2024), documents with dense layouts yield highly information-rich visual tokens. Compressing these risks severe information loss and degraded performance. Beyond vision, current methods also face inherent limits in other modalities. In automatic speech recognition (ASR) and automatic speech translation (AST) Ardila et al. (2019); Conneau et al. (2023), audio is encoded and then decoded into text using an MLLM Abouelenin et al. (2025); Chu et al. (2024). Audio tokens are dense and temporally continuous; pruning or merging them disrupts this continuity, leading to fragmented recognition or translation. Similarly, cross-lingual text translation may suffer significant degradation under high compression ratios.
4.2 Suboptimal Data Representation